# Interacting binaries on the Main Sequence as in-situ tracers of mass transfer efficiency and stability

### v1

@@ 3.3 Population properties
3.3.1 Rotationally limited mass accretion scheme
Figure 2 shows the allowed range of mass transfer efficiencies as a function of the current orbital period of each massive Algol. We observe that the current population of massive Algols does not require mass transfer to be very efficient at the shortest periods, unlike what is theoretically expected for rotationally limited accretion in the presence of strong tides. We observe evidence for an increase in the upper limit to the mass transfer efficiency as a function of the binary orbital period for systems in the Milky Way and the LMC. However, the mass transfer efficiency of the SMC systems shows no correlation with the orbital period (c.f. S. E. de Mink et al., 2007). Detailed binary evolution models predict that the majority of massive Algols at the shortest orbital period undergo fairly conservative mass transfer. However, the observed population of massive Algols do not populate the highest probability regions marked by the yellow background area in the $P_{\rm orb}-\varepsilon$ plane.
Comparing the top and middle panels of Fig. 3 in K. Sen et al. (2022), we see that most models that undergo highly efficient thermal timescale mass transfer ($\varepsilon>0.8$) merge during the Main Sequence after entering into a contact configuration. Most of the predicted population of massive Algols that do not merge on the Main Sequence undergoes inefficient mass transfer ($\varepsilon<0.5$). Therefore, the observed population of massive Algols may only represent the predicted population of Algols that do not merge on the Main Sequence. We speculate that highly efficient mass transfer at the shortest periods may lead to a contact phase on the Main Sequence (due to radial expansion of the accretor) at a much faster timescale than predicted by current binary evolution models (see also A. Menon et al., 2021; J. Henneco et al., 2024).
In the orbital period range 3-5 d, the observed systems in the Milky Way and LMC form two separate sub-populations, one requiring $\varepsilon<0.5$ and vice versa. In detailed binary evolution models, systems with initial mass ratios above 0.8 undergo fairly efficient mass transfer and vice versa (Fig. F2 of K. Sen et al., 2022). Therefore, the observed sub-populations of Algols in the Milky Way and the LMC may suggest a mass ratio dependence in mass transfer efficiency, which could also be an outcome of thermal timescale-limited mass transfer efficiency (e.g. K. Belczynski et al., 2008; F. R. N. Schneider et al., 2015).
@@ 3.3.1 Rotationally limited mass accretion scheme
3.3.2 Possible disk solutions
@@
Most of the current population of observed massive Algols in the SMC has mass ratios less than 2, whereas half of the Algol population in the LMC and Milky Way has mass ratios above 2. This may be an observational selection effect where it is harder to detect unequal mass ratio binaries in the SMC than the Milky Way (note that there are only seven observed massive Algols in the LMC in the current sample, and four of them are from the recent multi-epoch high-resolution VFTS-TMBM survey). However, there is a system with mass ratio above 3, and it is one of the least massive systems in our sample (Table 1, second row). If not biased by observational selection effects, our results strongly suggest that there may be a metallicity dependence in the mass transfer efficiency. A future multi-epoch spectroscopic survey of the remaining eclipsing binary sample from T. J. Harries et al. (2003) and R. W. Hilditch et al. (2005), can extend the observed sample and shed light on the lack of observed unequal mass ratio binaries in the SMC.
4 Discussion
4.1 Mass transfer efficiency and stability
@@
In K. Sen et al. (2022), we found that the orbital period and mass ratio distribution of massive Algols in the LMC and Milky Way give evidence for efficient mass transfer in some systems, while others have undergone inefficient mass transfer. We also found that the least massive systems are not well reproduced by their accretor spin-dependent mass transfer prescription. In this work, we show that the mass transfer stability criterion that requires the combined luminosity of two stars to be able to drive the excess mass loss via winds during an inefficient mass transfer episode is too strict to provide acceptable solutions for mass transfer efficiency in the least massive Algols (Fig. 6). S. E. de Mink et al. (2007) reported a similar lack of acceptable solutions for the least massive Algols in the SMC (see their Table 3, $\chi^{2}>10$), which may again be due to the lack of models that underwent stable mass transfer at low initial mass ratios. Mass loss from the L2 Lagrangian point (W. Lu et al., 2023) may provide an additional mechanism to remove the excess mass. However, L2 mass loss inherently implies a high value for $\gamma$, which may be ruled out by the upper limit on the maximum orbital period for Case A mass transfer (see blue regions, or the lack thereof, in the top four rows of the Fig. 5 and Fig. 6).
@@
In the SMC, S. E. de Mink et al. (2007) found a weak correlation between the mass transfer efficiency and initial orbital period of their best-fit binary models to the observed population of massive Algols. In our work, we also see a weak dependence of the allowed range of mass transfer efficiency on the current orbital period of the same sample, where the upper limit to the mass transfer efficiency from the rotationally limited accretion model decreases for increasing orbital period (Fig. 2). The detailed binary model grids of S. E. de Mink et al. (2007); C. Wang et al. (2020); K. Sen et al. (2022) assume that the mass lost due to inefficient mass transfer carries away the specific orbital angular momentum of the accretor. If $\varepsilon$ is assumed to be constant during the entire mass transfer episode, purple curves in Fig. 1 represent the initial orbital periods for various mass transfer efficiencies. However, the mass transfer efficiency in the detailed binary evolution models is not constant during the mass transfer phase (e.g. M. Renzo & Y. Götberg, 2021).
@@
The observed population of short-period stripped star+OB star binaries is expected to be the evolved counterparts of massive Algol binaries (S. Wellstein et al., 2001; K. Sen et al., 2022). At the high masses, these constitute the Wolf-Rayet+O star (WR+O) binaries. Most of the WR+O binaries have been found to require highly inefficient mass transfer, both in the Milky Way (J. Petrovic et al., 2005; Y. Shao & X.-D. Li, 2016; M. Nuijten & G. Nelemans, 2025) and the SMC (C. Schürmann & N. Langer, 2024). This is in agreement with our finding that a large majority of the massive Algols may have undergone inefficient mass transfer. On the other hand, low- and intermediate-mass Algols require a wide range of efficiencies to explain the observed populations (C. A. Nelson & P. P. Eggleton, 2001; R. Deschamps et al., 2013, 2015; N. Mennekens & D. Vanbeveren, 2017; A. Dervişoǧlu et al., 2018; S. H. Negu & S. B. Tessema, 2018).
@@
Recent studies on Be+sdOB binaries found that they require mass transfer to be significantly more efficient than in the case of massive Algols and WR+O binaries (T. Lechien et al., 2025; Y. Bao et al., 2025, see also A. Schootemeijer et al., 2018; L. Wang et al., 2021; A. Picco et al., 2025). Most of these systems are predicted to have undergone mass transfer after the main-sequence evolution of the donor star (Fig. 4 of T. Lechien et al., 2025), that is, they are not the direct evolutionary counterparts to Algol binaries. These systems strongly disfavour the accretion spin-up dependent mass transfer efficiency model (see also, A. Dervişoǧlu et al., 2010), as tides in long-period systems are less effective in tidally locking the accretor to the orbital angular velocity of the binary. The presence of a viscous decretion disk around the Be star in these binaries may allow for the loss of angular momentum via the boundary layer coupling between the star and the disk while continually accreting mass (B. Paczynski, 1991; R. Popham & R. Narayan, 1991; M. Colpi et al., 1991; G. S. Bisnovatyi-Kogan, 1993; S. E. de Mink et al., 2013; R. G. Martin et al., 2025).
@@
To reduce the predicted binary black hole merger rate in gravitational wave population synthesis models, I. Romero-Shaw et al. (2023) proposed conservative Case B mass transfer to increase the orbital periods of the resulting compact object binaries such that they do not merge within Hubble time. Similarly, previous studies propose that the mass transfer efficiency should be as high as 50% to explain the observed population of long-period Be X-ray binaries (Y. Shao & X.-D. Li, 2014; S. Vinciguerra et al., 2020; C. Schürmann et al., 2025), where such binaries are expected to have undergone Case B mass transfer. On the other hand, the observed population of short-period WR+O binaries in the local Universe seems to have undergone relatively inefficient mass transfer (J. Petrovic et al., 2005; Y. Shao & X.-D. Li, 2016; M. Nuijten & G. Nelemans, 2025). Likewise, M. Y. M. Lau et al. (2024) discuss that the observed population of gravitational wave sources may have undergone very inefficient mass transfer (see also, Y. Bouffanais et al., 2021; L. A. C. van Son et al., 2022; A. Picco et al., 2024). Recently, E. Zapartas et al. (2025) interpreted that the current sample of binary companions to stripped-envelope supernovae also favours inefficient mass accretion.
@@
C. Schürmann & N. Langer (2024, Fig. 8) showed that the lowest mass ratio for which L2 Overflow is avoided is a function of the mass transfer efficiency, and it is always greater than 0.5 for conservative mass transfer. While their models are at SMC metallicity, the accretors swell up even more at higher metallicities, and the parameter space for contact formation and L2 overflow increases to even higher initial mass ratios. This is consistent with our finding that the least massive Algols require inefficient mass transfer.
4.2 Extent of envelope stripping
@@
Figure 4: 2D-histogram showing the distribution of the donor masses as a function of the logarithm of initial donor mass during the Algol phase, for the detailed binary evolution models with LMC metallicity (K. Sen et al., 2022). The total probability is normalised such that the sum over all pixels for each logarithmic initial mass bin is unity (overplotted numbers show contribution from each pixel up to two significant digits). The red and blue curves show the convective core mass of the donor at Zero-Age Main Sequence with and without overshooting, respectively. The orange curve indicates that 80% of donor models during the Algol phase are stripped below the curve.
@@
Figure 4 shows the distribution of donor masses during the Algol phase. During the thermal timescale mass transfer on the Main Sequence, the donor may get stripped up to its initial convective core mass, which includes the overshooting region (e.g. Fig. 2 of C. Schürmann et al., 2024). The above is represented by the red curve in the 10-40 $M_{\odot}$ range of initial donor masses, where the majority of the donor masses at each initial donor mass bin reside during the Algol phase. During the nuclear timescale mass transfer phase, the donors lose further mass. We observe that the blue curve, representing the donor’s initial convective core mass excluding the overshooting region, provides an adequate approximation to the deepest extent of envelope stripping due to mass transfer on the Main Sequence (including both fast- and slow-Case A, e.g., J. Klencki et al. 2025).
@@
Physically, the response of the (inner layers of the) donor to the mass lost is primarily determined by its thermal state, that is the temperature gradient. Therefore, whether to include or not the convective boundary mixing region should be decided based on its gradient: for convective penetration (e.g., E. H. Anders et al., 2022), the gradient is adiabatic and no distinction with the convective core should be expected. However, other processes are known to contribute to convective boundary mixing, and the thermal gradient they establish on timescales that are long compared to the convective turnover is not always clear. In our study, we used the red curve (including convective boundary mixing) to determine the initial donor masses of the observed Algols (Sect. 2) as our fiducial extent of envelope stripping during the fast Case A mass transfer. Assuming a deeper extent of envelope stripping to estimate the initial donor mass in our analysis would lead to the inference of more inefficient mass transfer occurring in the observed systems.
@@
Finally, the extent of envelope stripping may non-trivially depend on the mass transfer efficiency itself, where the mass and angular momentum exchange between the donor and the accretor determines the closest orbit during the thermal timescale of mass transfer. The mass transfer efficiency in our models is dependent on both the initial orbital period and the initial mass ratio for a given initial donor mass. The detailed models cover the full range of mass transfer efficiencies (Fig. F2 of K. Sen et al., 2022). While quantifying the effect of mass transfer efficiency on the extent of envelope stripping is beyond the scope of this work, we note that the occurrence of both extremes in mass transfer efficiencies in our models also implies that the predicted spread of donor masses during the Algol phase in Fig. 4 includes the effect of mass transfer efficiency.
@@
Our detailed binary evolution models suggest that slow Case A mass transfer, occurring at the nuclear timescale, is highly efficient except at the longest initial orbital periods ($\varepsilon>0.6$, see Fig. F2 of K. Sen et al., 2022). The estimates for mass transfer efficiency derived in this work give the time-averaged mass transfer efficiency for each system (O. R. Pols & M. Marinus, 1994). Starting from the same initial donor mass, if the current donor mass is an outcome of mass lost during both fast and slow Case A, we expect that the prior thermal timescale mass transfer phase may have been slightly less efficient than our derived range of values for each observed massive Algol.
4.3 Convective core mass
@@
We calculate the initial mass of the donor from the initial mass of the convective core. The Ledoux criterion sets the initial convective core mass in our models for convection. Assuming the Schwarzschild criterion (e.g. in S. E. de Mink et al., 2007) for convection will lead to higher estimates of initial convective core masses for the same initial mass. Recent studies of gravity-mode asteroseismology of B stars (M. G. Pedersen et al., 2021) and analyses of eclipsing binaries (A. Claret & G. Torres, 2019; A. Tkachenko et al., 2020) suggest that the convective cores of stars are larger than assumed in many stellar evolution models. Building on this observational evidence, theoretical work has demonstrated that convective penetration in O- and B-type stars can enlarge the convective core by about 10–30% of the pressure scale height at the core boundary (E. H. Anders et al., 2022; A. S. Jermyn et al., 2022; C. Johnston et al., 2024). Larger convective cores in donors imply smaller envelopes for the same total initial mass. For systems where the current accretor mass is larger than the initial donor mass, the lower limit on mass transfer efficiency shifts to higher values. The initial mass ratio required for conservative mass transfer also relaxes to higher values.
@@
The detailed binary evolution models treat convective boundary mixing with parametric algorithms, which assume a radiative gradient in this region. The models of C. Wang et al. (2020); K. Sen et al. (2022) have a step overshooting parameter equal to 0.335 times the local pressure scale height and a radiative gradient in the overshooting region, calibrated to rotational velocity measurements from the FLAMES Survey of Massive Stars at $\sim$15 $M_{\odot}$ (I. Hunter et al., 2008; I. Brott et al., 2011, see also the recent theoretical study by R. Andrassy et al., 2024). N. Castro et al. (2014, 2018) later confirmed this calibration for stars of similar mass in the Milky Way and the SMC, but noted that stars of lower or higher mass appear to require correspondingly smaller or larger overshooting values (see also I. Baraffe et al., 2023; E. H. Anders & M. G. Pedersen, 2023; C. Johnston et al., 2024). Moreover, C. Johnston (2021) also find evidence of scatter in the convective core masses for the same total mass of an observed star, up to stellar masses $\leq$24 $M_{\odot}$. This scatter in convective core masses for the same total stellar mass may also be responsible for the separate sub-populations of massive Algols in the LMC and the Milky Way, which require either efficient or inefficient mass transfer at the same orbital period.
4.4 Wind mass loss

### v2

@@ III.3 Population properties
III.3.1 Rotationally limited mass accretion scheme
Figure 2 shows the allowed range of mass transfer efficiencies as a function of the current orbital period of each massive Algol. We observe that the current population of massive Algols does not require mass transfer to be very efficient at the shortest periods, unlike what is theoretically expected for rotationally limited accretion in the presence of strong tides. We observe evidence for an increase in the upper limit to the mass transfer efficiency as a function of the binary orbital period for systems in the Milky Way and the LMC. However, the mass transfer efficiency of the SMC systems shows no correlation with the orbital period (c.f. de Mink et al., 2007). Detailed binary evolution models predict that the majority of massive Algols at the shortest orbital period undergo fairly conservative mass transfer. However, the observed population of massive Algols do not populate the highest probability regions marked by the yellow background area in the $P_{\rm orb}-\varepsilon$ plane.
Comparing the top and middle panels of Fig. 3 in Sen et al. (2022), we see that most models that undergo highly efficient thermal timescale mass transfer ($\varepsilon>0.8$) merge during the Main Sequence after entering into a contact configuration. Most of the predicted population of massive Algols that do not merge on the Main Sequence undergoes inefficient mass transfer ($\varepsilon<0.5$). Therefore, the observed population of massive Algols may only represent the predicted population of Algols that do not merge on the Main Sequence. We speculate that highly efficient mass transfer at the shortest periods may lead to a contact phase on the Main Sequence (due to radial expansion of the accretor) at a much faster timescale than predicted by current binary evolution models (see also Menon et al., 2021; Henneco et al., 2024).
In the orbital period range 3-5 d, the observed systems in the Milky Way and LMC form two separate sub-populations, one requiring $\varepsilon<0.5$ and vice versa. In detailed binary evolution models, systems with initial mass ratios above 0.8 undergo fairly efficient mass transfer and vice versa (Fig. F2 of Sen et al., 2022). Therefore, the observed sub-populations of Algols in the Milky Way and the LMC may suggest a mass ratio dependence in mass transfer efficiency, which could also be an outcome of thermal timescale-limited mass transfer efficiency (e.g. Belczynski et al., 2008; Schneider et al., 2015).
@@ III.3.1 Rotationally limited mass accretion scheme
III.3.2 Possible disk solutions
@@
Most of the current population of observed massive Algols in the SMC has mass ratios less than 2, whereas half of the Algol population in the LMC and Milky Way has mass ratios above 2. This may be an observational selection effect where it is harder to detect unequal mass ratio binaries in the SMC than the Milky Way (note that there are only seven observed massive Algols in the LMC in the current sample, and four of them are from the recent multi-epoch high-resolution VFTS-TMBM survey). However, there is a system with mass ratio above 3, and it is one of the least massive systems in our sample (Table 2, second row). If not biased by observational selection effects, our results strongly suggest that there may be a metallicity dependence in the mass transfer efficiency. A future multi-epoch spectroscopic survey of the remaining eclipsing binary sample from Harries et al. (2003) and Hilditch et al. (2005), can extend the observed sample and shed light on the lack of observed unequal mass ratio binaries in the SMC.
IV Discussion
IV.1 Mass transfer efficiency and stability
@@
In Sen et al. (2022), we found that the orbital period and mass ratio distribution of massive Algols in the LMC and Milky Way give evidence for efficient mass transfer in some systems, while others have undergone inefficient mass transfer. We also found that the least massive systems are not well reproduced by their accretor spin-dependent mass transfer prescription. In this work, we show that the mass transfer stability criterion that requires the combined luminosity of two stars to be able to drive the excess mass loss via winds during an inefficient mass transfer episode is too strict to provide acceptable solutions for mass transfer efficiency in the least massive Algols (Fig. 6). de Mink et al. (2007) reported a similar lack of acceptable solutions for the least massive Algols in the SMC (see their Table 3, $\chi^{2}>10$), which may again be due to the lack of models that underwent stable mass transfer at low initial mass ratios. Mass loss from the L2 Lagrangian point (Lu et al., 2023) may provide an additional mechanism to remove the excess mass. However, L2 mass loss inherently implies a high value for $\gamma$, which may be ruled out by the upper limit on the maximum orbital period for Case A mass transfer (see blue regions, or the lack thereof, in the top four rows of the Fig. 5 and Fig. 6).
@@
In the SMC, de Mink et al. (2007) found a weak correlation between the mass transfer efficiency and initial orbital period of their best-fit binary models to the observed population of massive Algols. In our work, we also see a weak dependence of the allowed range of mass transfer efficiency on the current orbital period of the same sample, where the upper limit to the mass transfer efficiency from the rotationally limited accretion model decreases for increasing orbital period (Fig. 2). The detailed binary model grids of de Mink et al. (2007); Wang et al. (2020); Sen et al. (2022) assume that the mass lost due to inefficient mass transfer carries away the specific orbital angular momentum of the accretor. If $\varepsilon$ is assumed to be constant during the entire mass transfer episode, purple curves in Fig. 1 represent the initial orbital periods for various mass transfer efficiencies. However, the mass transfer efficiency in the detailed binary evolution models is not constant during the mass transfer phase (e.g. Renzo and Götberg, 2021).
@@
The observed population of short-period stripped star+OB star binaries is expected to be the evolved counterparts of massive Algol binaries (Wellstein et al., 2001; Sen et al., 2022). At the high masses, these constitute the Wolf-Rayet+O star (WR+O) binaries. Most of the WR+O binaries have been found to require highly inefficient mass transfer, both in the Milky Way (Petrovic et al., 2005; Shao and Li, 2016; Nuijten and Nelemans, 2025) and the SMC (Schürmann and Langer, 2024). This is in agreement with our finding that a large majority of the massive Algols may have undergone inefficient mass transfer. On the other hand, low- and intermediate-mass Algols require a wide range of efficiencies to explain the observed populations (Nelson and Eggleton, 2001; van Rensbergen et al., 2006, 2010; Deschamps et al., 2013, 2015; Mennekens and Vanbeveren, 2017; Dervişoǧlu et al., 2018; Negu and Tessema, 2018; van Rensbergen and de Greve, 2021). Evidence for a mass dependence of the accretion efficiency just in the massive star regime is discussed in Schürmann et al. (2025).
@@
Recent studies on Be+sdOB binaries found that they require mass transfer to be significantly more efficient than in the case of massive Algols and WR+O binaries (Lechien et al., 2025; Bao et al., 2025, see also Schootemeijer et al., 2018; Wang et al., 2021; Picco et al., 2025). Most of these systems are predicted to have undergone mass transfer after the main-sequence evolution of the donor star (Fig. 4 of Lechien et al., 2025), that is, they are not the direct evolutionary counterparts to Algol binaries. These systems strongly disfavour the accretion spin-up dependent mass transfer efficiency model (see also, Dervişoǧlu et al., 2010), as tides in long-period systems are less effective in tidally locking the accretor to the orbital angular velocity of the binary. The presence of a viscous decretion disk around the Be star in these binaries may allow for the loss of angular momentum via the boundary layer coupling between the star and the disk while continually accreting mass (Paczynski, 1991; Popham and Narayan, 1991; Colpi et al., 1991; Bisnovatyi-Kogan, 1993; de Mink et al., 2013; Martin et al., 2025).
@@
To reduce the predicted binary black hole merger rate in gravitational wave population synthesis models, Romero-Shaw et al. (2023) proposed conservative Case B mass transfer to increase the orbital periods of the resulting compact object binaries such that they do not merge within Hubble time. Similarly, previous studies propose that the mass transfer efficiency should be as high as 50% to explain the observed population of long-period Be X-ray binaries (Shao and Li, 2014; Vinciguerra et al., 2020; Schürmann et al., 2025), where such binaries are expected to have undergone Case B mass transfer. On the other hand, the observed population of short-period WR+O binaries in the local Universe seems to have undergone relatively inefficient mass transfer (Petrovic et al., 2005; Shao and Li, 2016; Nuijten and Nelemans, 2025). Likewise, Lau et al. (2024) discuss that the observed population of gravitational wave sources may have undergone very inefficient mass transfer (see also, Bouffanais et al., 2021; van Son et al., 2022; Picco et al., 2024). Recently, Zapartas et al. (2025) interpreted that the current sample of binary companions to stripped-envelope supernovae also favours inefficient mass accretion.
@@
Schürmann and Langer (2024, Fig. 8) showed that the lowest mass ratio for which L2 Overflow is avoided is a function of the mass transfer efficiency, and it is always greater than 0.5 for conservative mass transfer. While their models are at SMC metallicity, the accretors swell up even more at higher metallicities, and the parameter space for contact formation and L2 overflow increases to even higher initial mass ratios. This is consistent with our finding that the least massive Algols require inefficient mass transfer.
IV.2 Extent of envelope stripping
@@
Figure 4: 2D-histogram showing the distribution of the donor masses as a function of the logarithm of initial donor mass during the Algol phase, for the detailed binary evolution models with LMC metallicity (Sen et al., 2022). The total probability is normalised such that the sum over all pixels for each logarithmic initial mass bin is unity (overplotted numbers show contribution from each pixel up to two significant digits). The red and blue curves show the convective core mass of the donor at Zero-Age Main Sequence with and without overshooting, respectively. The orange curve indicates that 80% of donor models during the Algol phase are stripped below the curve.
@@
Figure 4 shows the distribution of donor masses during the Algol phase. During the thermal timescale mass transfer on the Main Sequence, the donor may get stripped up to its initial convective core mass, which includes the overshooting region (e.g. Fig. 2 of Schürmann et al., 2024). The above is represented by the red curve in the 10-40 $M_{\odot}$ range of initial donor masses, where the majority of the donor masses at each initial donor mass bin reside during the Algol phase. During the nuclear timescale mass transfer phase, the donors lose further mass. We observe that the blue curve, representing the donor’s initial convective core mass excluding the overshooting region, provides an adequate approximation to the deepest extent of envelope stripping due to mass transfer on the Main Sequence (including both fast- and slow-Case A, e.g., Klencki et al. 2025).
@@
Physically, the response of the (inner layers of the) donor to the mass lost is primarily determined by its thermal state, that is the temperature gradient. Therefore, whether to include or not the convective boundary mixing region should be decided based on its gradient: for convective penetration (e.g., Anders et al., 2022), the gradient is adiabatic and no distinction with the convective core should be expected. However, other processes are known to contribute to convective boundary mixing, and the thermal gradient they establish on timescales that are long compared to the convective turnover is not always clear. In our study, we used the red curve (including convective boundary mixing) to determine the initial donor masses of the observed Algols (Sect. II) as our fiducial extent of envelope stripping during the fast Case A mass transfer. Assuming a deeper extent of envelope stripping to estimate the initial donor mass in our analysis would lead to the inference of more inefficient mass transfer occurring in the observed systems.
@@
Finally, the extent of envelope stripping may non-trivially depend on the mass transfer efficiency itself, where the mass and angular momentum exchange between the donor and the accretor determines the closest orbit during the thermal timescale of mass transfer. The mass transfer efficiency in our models is dependent on both the initial orbital period and the initial mass ratio for a given initial donor mass. The detailed models cover the full range of mass transfer efficiencies (Fig. F2 of Sen et al., 2022). While quantifying the effect of mass transfer efficiency on the extent of envelope stripping is beyond the scope of this work, we note that the occurrence of both extremes in mass transfer efficiencies in our models also implies that the predicted spread of donor masses during the Algol phase in Fig. 4 includes the effect of mass transfer efficiency.
@@
Our detailed binary evolution models suggest that slow Case A mass transfer, occurring at the nuclear timescale, is highly efficient except at the longest initial orbital periods ($\varepsilon>0.6$, see Fig. F2 of Sen et al., 2022). The estimates for mass transfer efficiency derived in this work give the time-averaged mass transfer efficiency for each system (Pols and Marinus, 1994). Starting from the same initial donor mass, if the current donor mass is an outcome of mass lost during both fast and slow Case A, we expect that the prior thermal timescale mass transfer phase may have been slightly less efficient than our derived range of values for each observed massive Algol.
IV.3 Convective core mass
@@
We calculate the initial mass of the donor from the initial mass of the convective core. The Ledoux criterion sets the initial convective core mass in our models for convection. Assuming the Schwarzschild criterion (e.g. in de Mink et al., 2007) for convection will lead to higher estimates of initial convective core masses for the same initial mass. Recent studies of gravity-mode asteroseismology of B stars (Pedersen et al., 2021) and analyses of eclipsing binaries (Claret and Torres, 2019; Tkachenko et al., 2020) suggest that the convective cores of stars are larger than assumed in many stellar evolution models. Building on this observational evidence, theoretical work has demonstrated that convective penetration in O- and B-type stars can enlarge the convective core by about 10–30% of the pressure scale height at the core boundary (Anders et al., 2022; Jermyn et al., 2022; Johnston et al., 2024). Larger convective cores in donors imply smaller envelopes for the same total initial mass. For systems where the current accretor mass is larger than the initial donor mass, the lower limit on mass transfer efficiency shifts to higher values. The initial mass ratio required for conservative mass transfer also relaxes to higher values.
@@
The detailed binary evolution models treat convective boundary mixing with parametric algorithms, which assume a radiative gradient in this region. The models of Wang et al. (2020); Sen et al. (2022) have a step overshooting parameter equal to 0.335 times the local pressure scale height and a radiative gradient in the overshooting region, calibrated to rotational velocity measurements from the FLAMES Survey of Massive Stars at $\sim$15 $M_{\odot}$ (Hunter et al., 2008; Brott et al., 2011, see also the recent theoretical study by Andrassy et al., 2024). Castro et al. (2014, 2018) later confirmed this calibration for stars of similar mass in the Milky Way and the SMC, but noted that stars of lower or higher mass appear to require correspondingly smaller or larger overshooting values (see also Baraffe et al., 2023; Anders and Pedersen, 2023; Johnston et al., 2024). Moreover, Johnston (2021) also find evidence of scatter in the convective core masses for the same total mass of an observed star, up to stellar masses $\leq$24 $M_{\odot}$. This scatter in convective core masses for the same total stellar mass may also be responsible for the separate sub-populations of massive Algols in the LMC and the Milky Way, which require either efficient or inefficient mass transfer at the same orbital period.
IV.4 Wind mass loss